# Redistribution Through

### v2

@@ 3.4 Monotone segmentations and progressive pricing
The simple structure of strongly monotone segmentations and their uniqueness pave the way for a constructive approach. We describe a procedure, called the greedy redistributive algorithm, that outputs, for any market $\mu$, the unique strongly monotone segmentation of $\mu$ that is $\succsim_{\mathcal{R}}$-maximal. We explain in words the functioning of the algorithm and Appendix F contains its pseudo-code.
Strongly monotone segmentations can be seen as the result of lexicographic optimization: it benefits each type sequentially as much as possible, from the smallest to the largest. The greedy redistributive algorithm follows this logic and sequentially assigns types into segments. It starts by creating a lower segment with all buyers of type $\theta_{1}$ and then proceeds to include types $\theta_{2}$ in that segment. If all buyers $\theta_{2}$ can be included while keeping the price in the segment at $\theta_{1}$, the algorithm does so and proceeds to include types $\theta_{3}$. At some point in this process, the algorithm reaches a point such that including all buyers of a given type $\theta_{k}$ would violate Ob in that segment. When the algorithm reaches this point, it fits as many types $\theta_{k}$ as possible in the segment, so that Ob becomes binding at type $\theta_{k}$, and moves on to creating a new segment with the remaining mass of types $\theta_{k}$. This inclusion process is then repeated in the newly created segment starting at type $\theta_{k+1}$, and then proceeds sequentially as previously explained. The algorithm ends when every buyer has been assigned to some segment.
Link to optimal transport
Observe that the optimization problem $\mathrm{P}^{\star}_{w,\mu}$ is similar to the discrete optimal assignment problem (see Galichon, 2018, Chapter 3). The key difference is that instead of optimizing on couplings—the bivariate joint distributions with two prescribed marginals—we optimize on segmentations, which consist of joint distributions such that only the first marginal (the aggregate market $\mu$) is prescribed, while the second marginal (the price distribution $\sigma_{P}$) is endogenous and determined by the seller’s incentive constraints.
It is well established in statistics and optimal transport theory that strongly monotone joint distributions—also known as comonotone distributions—maximize the expectation of supermodular functions over the set of couplings (see, e.g., Theorem 2.1 in Puccetti and Wang, 2015, and Theorem 4.3 in Galichon, 2018). Therefore, since redistributive welfare functions are supermodular, it is natural that redistributive segmentations induce a positive-assortative matching of willingness to pay and prices. Moreover, just as weakly monotone segmentations correspond to the maximal elements of the redistributive order, comonotone distributions correspond to the maximal elements of the supermodular stochastic order (see claim (c) of Theorem 2.1 in Puccetti and Wang, 2015). [^24] As a side note, redistributive transfers (Definition 3) correspond to the mass transfers used in optimal transport to show that comonotone distributions are the extreme points of the transportation polytope (see Peyré and Cuturi, 2019, Proposition 3.4)—i.e. the set of discrete couplings.
Taken together, Propositions 1 and 2 and Theorems 2.A and 2.B thus provide a natural analogue in terms of segmentations to the characterizations of optimal assignments in optimal transport. However, since segmentations are constrained by the seller’s incentives, supermodularity alone no longer ensures comonotonicity. As discussed earlier, our concept of “strong-supermodularity” (Definition 6) becomes a necessary and sufficient condition for achieving comonotone optimal solutions.
3.5 Redistributive rents
@@
For any $\mu\in\Delta(\Theta)$, we say that a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ gives a rent to the seller if the profit under $\sigma$ is strictly higher than the profit under uniform pricing: that is, if
@@
$\sigma^{\star}_{\mu}(\theta,p)=\left\{\begin{array}[]{ll}\mu(\theta)&\text{if $\theta<p^{\star}_{\mu}$ and $p=\theta_{1}$}\\[15.0pt] \min\left\{\mu(p^{\star}_{\mu}),\displaystyle\frac{\theta_{1}}{p^{\star}_{\mu}-\theta_{1}}\displaystyle\sum_{\theta<p^{\star}_{\mu}}\mu(\theta)\right\}&\text{if $(\theta,p)=(p^{\star}_{\mu},\theta_{1})$}\\[15.0pt] \mu(p^{\star}_{\mu})-\sigma^{\star}_{\mu}(p^{\star}_{\mu},\theta_{1})&\text{if$(\theta,p)=(p^{\star}_{\mu},p^{\star}_{\mu})$}\\[15.0pt] \mu(\theta)&\text{if $\theta>p^{\star}_{\mu}\text{ and }p=p^{\star}_{\mu}$}\\[15.0pt] 0&\text{otherwise.}\end{array}\right.$
@@
By construction, $\sigma^{\star}_{\mu}$ satisfies $\mathrm{C}_{\mu}$ and $\operatorname{supp}(\sigma^{\star}_{\mu})\subset\Omega$. Therefore, if $\sigma^{\star}_{\mu}$ satisfies Ob then it is a segmentation belonging to $\Sigma_{\Omega}(\mu)$. In this case, we refer to $\sigma^{\star}_{\mu}$ as being feasible. Note that, if $\sigma^{\star}_{\mu}$ is feasible, then it is the unique $\succsim_{\mathcal{R}}$-maximal and strongly monotone segmentation of [^25] Note that $\sigma^{\star}_{\mu}$ is constructed such that, if it is feasible, the seller is indifferent between charging $\theta_{1}$ and $p^{\star}_{\mu}$ in segment $\theta_{1}$. As such, $\sigma^{\star}_{\mu}$ satisfies conditions (i) and (ii) of Proposition 2 if it is feasible, implying that it is $\succsim_{\mathcal{R}}$-maximal. $\mu$. Furthermore, $\sigma^{\star}_{\mu}$ has a very simple structure. It assigns all types lower than $p^{\star}_{\mu}$ to a “discount” segment where the lowest possible price $\theta_{1}$ is charged, and all the remaining types to a residual segment where the uniform price $p^{\star}_{\mu}$ is charged. Figure 6 illustrates $\sigma^{\star}_{\mu}$.
Theorem 3 (Redistributive Rents).
For any $\mu\in\Delta(\Theta)$ and any strongly redistributive welfare function $w$, the two following statements are true.
@@
A segmentation does not give any rent to the seller if and only if $p^{\star}_{\mu}$ is an optimal price in every market segment. [^26] This follows easily from the fact that the seller’s total profit is the sum of its profits on each segment. Therefore, a strongly monotone segmentation does not give a rent to the seller only if it satisfies two conditions. First, it must generate only two market segments. One containing every type weakly lower than $p^{\star}_{\mu}$, and one containing every type weakly greater than $p^{\star}_{\mu}$. Second, Ob must bind at $p^{\star}_{\mu}$ in the segment $\theta_{1}$. It can easily be seen that $\sigma^{\star}_{\mu}$ is the only strongly monotone segmentation that satisfies those two requirements. Furthermore, following Theorem 2.B, for any strongly redistributive $w$, the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is the unique strongly monotone one that is $\succsim_{\mathcal{R}}$-maximal. Therefore, if $\sigma^{\star}_{\mu}$ is feasible, then it is uniquely $\mathrm{P}^{\star}_{w,\mu}$-optimal and does not give a rent to the seller. On the contrary, if $\sigma^{\star}_{\mu}$ is not feasible, then the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation must give a rent to the seller. ∎
@@
Theorem 3 highlights the key tradeoff generated by redistributive concerns. Benefiting poor consumers may require imposing a sacrifice on richer consumers’ surplus. Strongly redistributive welfare functions precisely characterize when the welfare gains from redistribution more than compensate for the surplus loss of higher types. Given that redistributive segmentations are efficient (Lemma 1), as a byproduct of this loss, the seller benefits from a redistributive rent. This result is easily explainable in light of the structure of strongly redistributive segmentations (Theorem 2.B). Such segmentations pool as many low-types as possible in low-price segments, allowing them to benefit from lower prices. Consequently, high types are separated from low types, enabling the seller to better price discriminate. This increases the seller’s profit, except in the special case where $\sigma^{\star}_{\mu}$ is feasible.
4 Implementation through price regulation
@@
If the regulator can verify the segmentation chosen by the seller, such an implementation problem is trivial. The regulator can simply commit to punishing the seller if his chosen segmentation differs from her preferred one. In practice, however, it is unlikely that the regulator can perfectly observe the segmentation set by the seller. We show that observing the price distribution induced by a segmentation is sufficient for the regulator to implement her preferred segmentation.
Assume that the regulator and the seller both know the aggregate market $\mu\in\Delta(\Theta)$. The regulator communicates to the seller the segmentation $\sigma$ she wishes to implement, after which the seller chooses a market segmentation $\sigma^{\prime}\in\Sigma(\mu)$, which may or may not coincide with the regulator’s one. The regulator is unable to observe $\sigma^{\prime}$ directly but can verify and enforce punishments based on the induced price distribution [^27] As discussed in the introduction, observing the distribution of prices seems feasible; see Footnote 6. $\sigma^{\prime}_{P}$. Consequently, for any $\mu\in\Delta(\Theta)$, the set of undetectable deviations from $\sigma$ is
@@
A segmentation can be implemented through price-based regulation if there exists no strictly profitable deviation for the seller that is undetectable by the regulator. The next proposition shows that efficiency is, in fact, sufficient for implementability. [^28] Our notion of implementability is equivalent to the notion of credibility defined by Lin and Liu (2024) in the context of Bayesian persuasion. Specifically, the authors define an information structure as credible if there exists no alternative information structure that would strictly increase the sender’s payoff while unaffecting the marginal distribution of signal realizations. As such, Proposition 3 shows that any efficient segmentation is also seller-credible. Our notion of implementation is also equivalent to the notion of price-regulation defined by Schlom (2024) in the context of second-degree price discrimination.
@@
Let $\mu\in\Delta(\Theta)$. If $\sigma\in\Sigma_{\Omega}(\mu)$, then $\sigma$ is implementable through price-based regulation.
@@
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$. Recall that, for any $(\theta,p)\in\Omega$ we must have that $\sigma(\theta,p)=\sigma(\theta\,|\,p)\,\sigma_{P}(p)$. We thus obtain the following decomposition of the platform’s profit under $\sigma$:
$\sum_{p\in\operatorname{supp}(\sigma_{P})}p\sum_{\theta\geq p}\sigma(\theta,p)$ $=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p)\overbrace{\sum_{\theta\geq p}\sigma(\theta\,|\,p)}^{\mathclap{\text{$=1$ since $\sigma\in\Sigma_{\Omega}(\mu)$}}}$
@@
Therefore, any efficient segmentation and, following Lemma 1, any redistributive segmentation, can be implemented through price-based regulation. The proof follows from a simple observation. Namely, the difference in the seller’s profit between two segmentations that induce the same price distribution only depends on the mass of consumers that buy the good at each price. For efficient segmentations, this mass is maximal because every buyer consumes. Therefore, there cannot exist any strictly beneficial deviation.
5 Concluding remarks

### v5

@@ 3.3.2. Strengthening Monotonicity
The simple structure of strongly monotone segmentations and their uniqueness pave the way for a constructive approach. We describe a procedure, called the greedy redistributive algorithm, that outputs, for any market $\mu$, the unique strongly monotone segmentation of $\mu$ that is saturated. We explain in words the functioning of the algorithm and LABEL:secap:greedy contains its pseudo-code.
Strongly monotone segmentations can be seen as the result of lexicographic optimization: it benefits each type sequentially as much as possible, from the smallest to the largest. The greedy redistributive algorithm follows this logic and sequentially assigns types into segments. It starts by creating a lower segment with all buyers of type $\theta_{1}$ and then proceeds to include types $\theta_{2}$ in that segment. If all buyers $\theta_{2}$ can be included while keeping the price in the segment at $\theta_{1}$, the algorithm does so and proceeds to include types $\theta_{3}$. At some point in this process, the algorithm reaches a point such that including all buyers of a given type $\theta_{k}$ would violate Ob in that segment. When the algorithm reaches this point, it fits as many types $\theta_{k}$ as possible in the segment so that Ob becomes binding for some price and moves on to creating a new segment with the remaining mass of types $\theta_{k}$. This inclusion process is then repeated in the newly created segment starting at type $\theta_{k+1}$, and then proceeds sequentially as previously explained. The algorithm ends when every buyer has been assigned to some segment.
3.4.  Redistributive Rents
@@
For any $\mu\in\Delta(\Theta)$, we say that a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ gives a rent to the seller if the profit under $\sigma$ is strictly higher than the profit under uniform pricing: that is, if
@@
[…]
@@
By construction, $\sigma^{\star}_{\mu}$ satisfies $\mathrm{C}_{\mu}$ and $\operatorname{supp}(\sigma^{\star}_{\mu})\subset\Omega$. Therefore, if $\sigma^{\star}_{\mu}$ satisfies Ob then it is a segmentation belonging to $\Sigma^{\mathrm{eff}}(\mu)$. In this case, we refer to $\sigma^{\star}_{\mu}$ as being feasible. Note that, if $\sigma^{\star}_{\mu}$ is feasible, then it is the unique saturated and strongly monotone segmentation of [^30] Note that $\sigma^{\star}_{\mu}$ is constructed such that, if it is feasible, the seller is indifferent between charging $\theta_{1}$ and $p^{\star}_{\mu}$ in segment $\theta_{1}$. As such, $\sigma^{\star}_{\mu}$ satisfies conditions (a) and (b) of Definition 1 if it is feasible, implying that it is saturated. Intuitively, the feasibility of $\sigma^{\star}_{\mu}$ is therefore demanding as it requires a large mass of type $p^{\star}_{\mu}$ such that $p^{\star}_{\mu}$ is an optimal price in both segments. $\mu$. Furthermore, $\sigma^{\star}_{\mu}$ has a very simple structure. It assigns all types lower than $p^{\star}_{\mu}$ to a “discount” segment where the lowest possible price $\theta_{1}$ is charged, and all the remaining types to a residual segment where the uniform price $p^{\star}_{\mu}$ is charged. Figure˜6 illustrates $\sigma^{\star}_{\mu}$.
Proposition 3 (Redistributive Rents).
For any $\mu\in\Delta(\Theta)$ and any strongly redistributive welfare function $w$, the two following statements are true:
@@
A segmentation does not give any rent to the seller if and only if $p^{\star}_{\mu}$ is an optimal price in every market segment. [^31] This follows easily from the fact that the seller’s total profit is the sum of its profits on each segment. Therefore, a strongly monotone segmentation does not give a rent to the seller only if it satisfies two conditions. First, it must generate only two market segments. One containing every type weakly lower than $p^{\star}_{\mu}$, and one containing every type weakly greater than $p^{\star}_{\mu}$. Second, Ob must bind at $p^{\star}_{\mu}$ in the segment $\theta_{1}$. It can easily be seen that $\sigma^{\star}_{\mu}$ is the only strongly monotone segmentation that satisfies those two requirements. Furthermore, following Proposition˜2, for any strongly redistributive $w$, the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is the unique strongly monotone one that is saturated. Therefore, if $\sigma^{\star}_{\mu}$ is feasible, then it is uniquely $\mathrm{P}^{\star}_{w,\mu}$-optimal and does not give a rent to the seller. On the contrary, if $\sigma^{\star}_{\mu}$ is not feasible, then the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation must give a rent to the seller. ∎
@@
Proposition˜3 highlights the key tradeoff generated by redistributive concerns. Benefiting poor consumers may require imposing a sacrifice on richer consumers’ surplus. Strongly redistributive welfare functions precisely characterize when the welfare gains from redistribution more than compensate for the surplus loss of higher types. Given that redistributive segmentations are efficient (Lemma˜1), as a byproduct of this loss, the seller benefits from a redistributive rent. This result is easily explainable in light of the structure of strongly redistributive segmentations (Proposition˜2). Such segmentations pool as many low-types as possible in low-price segments, allowing them to benefit from lower prices. Consequently, high types are separated from low types, enabling the seller to better price discriminate. This increases the seller’s profit, except in the special case where $\sigma^{\star}_{\mu}$ is feasible.
3.5.  Discussion
3.5.1. Redistribution and Consumer-Surplus Maximization
The results in this paper highlight a variety of ways in which redistributive segmentations can differ from consumer surplus-maximizing segmentations (bbm15). Notably, redistributive segmentations might differ from consumer surplus-maximizing (henceforth CS-maximizing) ones in how they assign consumers to segments, in the prices they generate, and in the way they distribute surplus across the buyers and the seller.
Redistributive segmentations exhibit a positive-assortative structure, assigning consumers with a lower willingness to pay to segments with low prices and consumers with a higher willingness to pay to segments with higher prices. As seen in Section˜3.3, all redistributive segmentations satisfy weak monotonicity: if a buyer of type $\theta$ consumes at price $p$, there must exist a buyer of higher type who consumes at a higher price. When the redistributive motive is sufficiently strong, the optimal redistributive segmentations also satisfy strong monotonicity: if a buyer of type $\theta$ consumes at price $p$, every buyer of a higher type must be consuming at a higher price. CS-maximizing segmentations need not satisfy even this weak notion of monotonicity. For instance, bbm15’ (bbm15) greedy algorithm may not lead to a weakly monotone segmentation (taking the equivalent direct segmentation).
The monotone structure of redistributive segmentations qualifies a sense in which redistributive segmentations are “more separating” than CS-maximizing ones. This greater separation is echoed in the distributions of prices induced by such segmentations: while CS-maximizing segmentations only induce segments with discounts relative to the uniform price, redistributive segmentations might also induce segments with prices higher than the uniform. As such, prices might be more dispersed under redistributive segmentations. Another symptom of this greater separation is the distribution of surplus across sides of the market. As treated in Section˜3.4, while CS-maximizing segmentations allocate all of the surplus generated by the segmentation to buyers, redistributive segmentations might, as a byproduct of prioritizing the welfare of poorer consumers, grant extra surplus to the seller.
Finally, it is important to note that the set of redistributive segmentations intersects with the set of CS-maximizing ones. Indeed, as a welfare function, total consumer surplus is included in the class of welfare functions we study, despite not being our focus since condition R, which captures our notion of redistributive concerns, is satisfied everywhere with equality. Regardless, from Theorem˜1 it follows that in the set of CS-maximizing segmentations, one can always find a saturated one.
The redistributive order characterized in Proposition˜1 is also helpful if one wants to use redistribution as a criterion to select among CS-maximizing segmentations. CS-maximizing segmentations cannot be improved in the redistributive order through downward transfers, but might be improvable through redistributive transfers. Departing from some CS-maximizing segmentation, sequentially performing feasible redistributive transfers will eventually result in a saturated segmentation that is also CS-maximizing.
3.5.2. Link to Stochastic Orders and Optimal Transport
Our characterization of the redistributive order, as presented in Proposition˜1, conceptually parallels the characterization of the supermodular stochastic order developed independently by Muller2013 and Meyer2015. The critical distinction lies in their respective objects of comparison: the supermodular stochastic order compares couplings—bivariate joint distributions with two prescribed marginals—whereas the redistributive order compares segmentations, which consist of joint distributions where only the first marginal (the aggregate market $\mu$) is prescribed, while the second marginal (the price distribution $\sigma_{P}$) emerges endogenously from the seller’s incentive constraints. [^32] This observation is generally valid for any persuasion problem, where the prior distribution of the state is prescribed while the distribution over the receiver’s action is determined by the obedience constraints (see Bergemann2019). As such, our characterization could be fruitfully extended to answer more generally a recurring question in the literature: When are optimal information structures monotone?
Therefore, the characterization of the supermodular order over couplings differs from that of the redistributive order over segmentations in a substantive way. The supermodular order is characterized only in terms of supermodular transfers (akin to redistributive transfers in our terminology), whereas the redistributive order is characterized in terms of both supermodular transfers, which preserve both marginals, and downward transfers, which can modify the marginal distribution of prices. This distinction arises directly from the endogenous nature of segmentations’ second marginal. Since downward transfers inevitably alter the price distribution, they cannot be applied to couplings, where both marginals are prescribed. [^33] It is noteworthy that redistributive transfers (Definition 4) correspond to the mass transfers used in optimal transport to show that the comonotone distributions are the extreme points of the transport polytope (see Carlier2022, Theorem 6.14). It could be worthwhile for future research to use similar tools to characterize the extreme points of the “persuasion polytope.”
Observe also that the optimization problem $\mathrm{P}^{\star}_{w,\mu}$ bears similarity to the discrete optimal assignment problem (see Galichon2018, Chapter 3). Again, the key difference is that rather than maximizing the expectation of a function on the set of couplings, we maximize it on the set of (efficient) segmentations.
In statistics and optimal transport theory it is known that comonotone joint distributions maximize the expectation of supermodular functions on the set of couplings (see, e.g., Theorem 2.1 in Puccetti2015, and Theorem 4.3 in Galichon2018). We show that, similarly as for couplings, the supermodularity of redistributive welfare functions induces a positive-assortative matching of willingness to pay and prices. Furthermore, we show in Lemma˜2 that saturated segmentations correspond to the maximal elements of the redistributive order, which provides a parallel to how comonotone distributions correspond to the maximal elements of the supermodular stochastic order (see claim (c) of Theorem 2.1 in Puccetti2015).
Collectively, Theorem˜1, Corollaries˜1 and 2 thus provide a natural analog for segmentations to the characterizations of (discrete) optimal assignments in optimal transport theory. However, since segmentations are constrained by the seller’s incentives, supermodularity alone no longer guarantees comonotonicity. As discussed earlier, our concept of strongly redistributive welfare functions (Definition˜6), which can be understood as a form of “strong-supermodularity,” becomes both necessary and sufficient to achieve comonotone optimal solutions.
4. Implementation Through Price Regulation
@@
If the regulator can verify the segmentation chosen by the seller, this implementation problem is trivial. The regulator can simply commit to punishing the seller if his chosen segmentation differs from her preferred one. However, it is unlikely that the regulator can perfectly observe the segmentation set by the seller. We show that observing the price distribution induced by segmentations is sufficient for the regulator to implement her preferred segmentation.
Assume that both the regulator and the seller know the aggregate market $\mu\in\Delta(\Theta)$. The regulator communicates to the seller the segmentation $\sigma$ she wishes to implement, after which the seller chooses a market segmentation $\sigma^{\prime}\in\Sigma(\mu)$. The regulator cannot observe $\sigma^{\prime}$ directly but can verify and enforce punishments based on the induced price distribution [^34] As discussed in the introduction, observing the distribution of prices is realistic; see Footnote 7. $\sigma^{\prime}_{P}$. Consequently, for any $\mu\in\Delta(\Theta)$, the set of undetectable deviations from $\sigma$ is
@@
A segmentation can be implemented through price-based regulation if there exists no strictly profitable deviation for the seller that is undetectable by the regulator. The next proposition shows that efficiency is, in fact, sufficient for implementability. [^35] Our notion of implementability is equivalent to the notion of credibility defined by linliu in the context of Bayesian persuasion. Specifically, the authors define an information structure as credible if there exists no alternative information structure that would strictly increase the sender’s payoff while unaffecting the marginal distribution of signal realizations. As such, Proposition 4 shows that any efficient segmentation is also seller-credible. Our notion of implementation is also equivalent to the notion of price-regulation defined by Schlom2024 in the context of second-degree price discrimination.
@@
Let $\mu\in\Delta(\Theta)$. If $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$, then $\sigma$ is implementable through price-based regulation.
@@
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$. Recall that, for any $(\theta,p)\in\Omega$ we must have that $\sigma(\theta,p)=\sigma(\theta\,|\,p)\,\sigma_{P}(p)$. We thus obtain the following decomposition of the platform’s profit under $\sigma$:
$\sum_{p\in\operatorname{supp}(\sigma_{P})}p\sum_{\theta\geq p}\sigma(\theta,p)$ $=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p)\overbrace{\sum_{\theta\geq p}\sigma(\theta\,|\,p)}^{\mathclap{\text{$=1$ since $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$}}}$
@@
Therefore, any efficient segmentation and, following Lemma˜1, any redistributive segmentation, can be implemented through price-based regulation. The proof follows from a simple observation: the difference in the seller’s profit between two segmentations that induce the same price distribution only depends on the mass of consumers that buy the good at each price. For efficient segmentations, this mass is maximal because every buyer consumes. Therefore, there cannot be any strictly beneficial deviation.
5. Concluding Remarks